# references.bib

### v1

@@ 3.2 Generator mask
Zonotopes are a scalable convex set representation. By definition (see Section 2.2), they map a point from the hypercube in the generator space to a point in the set space. The generator mask is based on exploiting this property by letting the RL agent select actions in the generator space. Since the size of the output layer of the policy or Q-function network cannot change during the learning process, we require that the dimensionality of the generator space static. Overall, the generator mask requires the following assumption:
@@
The relevant action set $\mathcal{A}^{r}(s)$ is represented by a zonotope and the generator dimensions for the relevant action set are fixed during training and deployment.
In particular, the relevant action set $\mathcal{A}^{r}(s)$ is defined as $\langle c(s),G(s)\rangle_{\mathcal{Z}}$, with $G\in\mathbb{R}^{P\times N}$ and $c\in\mathbb{R}^{N}$. Note that in practice Assumption 3 is typically a weak limitation since a high-dimensional generator space allows for a greater variety of relevant action sets. Since the policy $\pi_{\theta}(a|s)$ is defined in the space of the zonotope’s hypercube, we treat the hypercube as a latent action space $\mathcal{A}^{l}\subset\mathbb{R}^{P}$. However, the domain of $\pi_{\theta}^{r}(a^{r}|s)$ is still a subset of the action space $\mathcal{A}^{r}\subseteq\mathcal{A}\subset\mathbb{R}^{N}$ (see Figure 1(b)).
For deriving the policy gradient of the generator mask, we assume that the policy follows the most common probability distribution for continuous RL policies:
@@
$\pi_{\theta}(a|s)$ follows a normal distribution $\mathcal{N}(a;\mu_{\theta},\Sigma_{\theta})$.
@@
The relevant policy of the generator mask is given by
$\pi_{\theta}^{r}(a|s)=\mathcal{N}(a;G\mu_{\theta}+c,G\Sigma_{\theta}G^{T})$ (11)
@@
The mapping function for the generator mask $g:\mathcal{A}\rightarrow\mathcal{A}^{r}$, as defined by the mapping of a zonotope,
$a^{r}=g(a)=c+Ga,$ (12)
represents a linear mapping of $a\sim\pi_{\theta}(a|s)$. Therefore, we can apply the linear transformation theorem for multivariate normal distributions, to compute the closed-form of the relevant policy in (11) (see Lemma 1 in Appendix A.1). ∎
Next, we discuss the implications of this mapping on the policy gradient. Since $g(a)$ is not bijective generally, we cannot derive the gradient through the change of variables formula as for the ray mask.
@@
The policy gradient for $\pi_{\theta}^{r}(a^{r}|s)$ as defined in (11) with respect to $\mu_{\theta}$ and $\Sigma_{\theta}$ is given by
@@
The proposition is proven with linear algebra operations and detailed in Appendix A.2. ∎
Note that the expressions in (13) and (14) simplify to $\nabla_{\mu_{\theta}}\log\pi_{\theta}(a|s)$, and $\nabla_{\Sigma\theta}\log\pi_{\theta}(a|s)$, respectively, if $G^{-1}$ exists, i.e., $G$ is square and non-singular (see Proposition 5 in Appendix A.2).
@@ 3.3 Distributional mask
The intuition behind the distributional mask is drawn from discrete action masking, where the probability for irrelevant actions is set to 00 [Huang2022]. For continuous action spaces, we aim to achieve the same by ensuring that actions are only sampled from the relevant action set $\mathcal{A}^{r}(s)$, by constraining the policy distribution $\pi_{\theta}(a|s)$ to this convex set (see Figure 1(c)). For the one-dimensional case, this can be expressed by the truncated distribution. In higher dimensions, the resulting policy density is
$\pi_{\theta}^{r}(a^{r}|s)=\frac{\phi(a,s)\pi_{\theta}(a|s)}{\int_{\mathcal{A}^{r}}\pi_{\theta}(a|s)\mathrm{d}a},$ (15)
@@
$\phi(a,s)=\begin{cases}1&\text{if }a\in\mathcal{A}^{r}(s),\\ 0&\text{otherwise}.\end{cases}$ (16)
Since there is no closed-form of this density, we employ Markov-Chain Monte-Carlo sampling to sample actions from the policy. More specifically, we utilize the random direction Hit-and-Run algorithm: a geometric random walk that allows sampling from a non-negative, integrable function $f:\mathbb{R}^{N}\rightarrow\mathbb{R}_{+}$, while constraining the samples to a bounded set [zabinsky_hit-and-run_2013]. The algorithm iteratively chooses a random direction from the current point, computes the one-dimensional, truncated probability density of $f$ along this direction, and samples a new point from this density. The approach is particularly effective for high-dimensional spaces where other sampling methods might struggle with convergence or efficiency. For the distributional mask, $f$ is the policy $\pi_{\theta}(a|s)$, and the set is $\mathcal{A}^{r}(s)$. As suggested by [lovasz_hit-and-run_2006], we choose the mixing time to be $N^{3}$, i.e., we execute $N^{3}$ iterations before accepting the sample. To estimate the integral in (15), we use numerical integration with cubature [genz_adaptive_2003], which is a method to approximate the definite integral of a function $h:\mathbb{R}^{N}\rightarrow\mathbb{R}$ over a multidimensional geometric set.
@@
$\nabla_{\theta}\log\pi_{\theta}^{r}(a^{r}|s)=\nabla_{\theta}\log\pi_{\theta}(a|s)-\nabla_{\theta}\log\int_{\mathcal{A}^{r}}\pi_{\theta}(a|s)\mathrm{d}a.$ (17)
@@
The indicator function, $\pi_{\theta}^{r}(a^{r}|s)$ is not continuous and differentiable, which necessitates the use of the sub-gradient for learning. However, since $a^{r}\in\mathcal{A}^{r}$ always holds, the gradient has to be computed for the continuous part of $\pi_{\theta}^{r}(a^{r}|s)$ only, and thus $\phi(a,s)$ can be omitted. ∎
Since we cannot compute the gradient of the numerical integral $\int_{\mathcal{A}^{r}}\pi_{\theta}(a|s)\mathrm{d}a$, we treat the integral as a constant in practice, and approximate $\nabla_{\theta}\log\pi_{\theta}^{r}(a^{r}|s)\approx\nabla_{\theta}\log\pi_{\theta}(a|s)$. We discuss potential improvements in Section 4.3.
@@ 4.1 Environments
We briefly introduce the three environments with their dynamics, state space, action space, reward, and relevant action set $\mathcal{A}^{r}(s)$. Parameters for the environments are detailed in Appendix A.5.
@@ 4.1.1 Seeker Reach-Avoid
This episodic environment features an agent navigating a 2D space, tasked with reaching a goal while avoiding a circular obstacle (see Figure 2). It is explicitly designed to provide an intuitive relevant action set $\mathcal{A}^{r}$. The system is defined as in Section 2.2: The dynamics for the agent position $s=\left[s_{x},s_{y}\right]$ and the action $a=\left[a_{x},a_{y}\right]$ are $\dot{s}=a$, and there are no disturbances.
$\mathcal{S}^{r}$$s^{*}$$\mathcal{O}$$\mathcal{S}_{k+1}$$s_{k}$$\mathcal{A}$$\mathcal{A}^{r}_{k}$
@@
Figure 2: The Seeker Reach-Avoid environment with examplatory state space configuration (left) and action space (middle). The agent (black) has to reach the goal (gray), while avoiding the obstacle (red). The center of the action space is illustrated with a black dot and the relevant action set $\mathcal{A}^{r}$ for the current state is shown in green. The state set at the next time step $k+1$ reachable with the relevant action set is $\mathcal{S}_{k+1}$.
The environment is characterized by the agent’s position, the goal position $s^{*}$, the obstacle position $o$, and the obstacle radius $r_{o}$. These values are pseudo-randomly assigned at the beginning of each episode, with the constraints that the goal cannot be inside the obstacle and the obstacle blocks the direct path between the agent’s initial position and the goal. The reward for each time step is
$r(a,s)=\begin{cases}100&\text{if goal reached},\\ -100&\text{if collision occurred},\\ -1-||s^{*}-s||_{2}&\text{otherwise}.\end{cases}$ (18)
We compute the relevant action set $\mathcal{A}^{r}(s)$ so that all actions that cause a collision with the obstacle or the boundary, are excluded (see Appendix A.3.3).
@@
The 2D Quadrotor environment models a stabilization task and employs an action space where the two action dimensions are coupled, i.e., rotational movement is originating from differences between the action values and vertical movement is proportional to the sum of the action values. Specifically, the system dynamics are:
where the state is $s=\left[x,z,\dot{x},\dot{z},\theta,\dot{\theta}\right]$, the action is $a=\left[a_{1},a_{2}\right]$, and the disturbance is $w=\left[w_{1},w_{2}\right]$. The parameter values are specified in Appendix A.5. The relevant action set is computed based on the system dynamics and a relevant state set (see Appendix, Eq. (35)). The reward function is defined as
$r(a,s)=\exp\left(-\|s-s^{*}\|_{2}-\frac{0.01}{2}\left\|\frac{a-\mathrm{min}(\mathcal{A})}{\mathrm{max}(\mathcal{A})-\mathrm{min}(\mathcal{A})}\right\|_{1}\right),$ (20)
where $s^{*}=\mathbf{0}$ is the stabilization goal state, $\mathrm{min}(\mathcal{A})$ returns the lower bound for each dimension of the action space, and $\mathrm{max}(\mathcal{A})$ the upper bound.
@@
The last environment models a stabilization task defined in [Kaynama2015]. The system is modeled in a twelve-dimensional state space with state $s=[x,y,z,\dot{x},\dot{y},\dot{z},\theta,\phi,\psi,\dot{\theta},\dot{\phi},\dot{\psi}]$ and four-dimensional action space with action $a=[a_{1},a_{2},a_{3},a_{4}]$. The system dynamics are $\dot{s}=[\dot{x},\dot{y},\dot{z},-9.81\phi,9.81\theta,a_{1},\dot{\theta},\dot{\phi},\dot{\psi},a_{2},a_{3},a_{4}]$ The reward is defined in (20) As for the 2D Quadrotor environment, we calculate the relevant action set based on the system dynamics and a relevant state set (see Appendix, Eq. (35)).
@@ 4.2 Results
The reported training results are based on ten random seeds for each configuration. Figure 3 shows the mean reward and bootstrapped 95% confidence interval [patterson2023empirical] for the three environments. For all environments, the baseline, which is standard PPO, converges significantly slower or not at all. Additionally, the initial rewards when using action masking methods are significantly higher than for the baseline, indicating that the exploration is indeed constrained to relevant task-fulfilling actions. We depict the mean and standard deviation of the episode return during deployment in Table 1. The baseline episode returns are significantly lower than for action masking while the three masking methods perform similar.
For the Seeker environment, the fraction between relevant action set and global action space is on average $70\%$. All action masking methods converge significantly faster to high rewards than the baseline. The generator mask and the distributional mask achieve the highest final reward. For the 2D Quadrotor, the relevant action to action space fraction is much smaller with on average $28\%$. In this environment, the ray mask and generator mask achieve the highest reward. While the baseline converges significantly slower, the final average reward is similar to the one of the distributional mask. Note that the confidence interval for the baseline is significantly larger. This is because three of the ten runs do not converge and produce an average reward around one throughout training. Note that if we constrain the relevant action set to an interval for this environment, our optimization problem in (35) often renders infeasible because the maximal relevant action set cannot be well approximated by an interval. Thus, the masking method proposed in [Krasowski2023b.ProvablySafeRLSurvey] is not suitable for the 2D Quadrotor task. The results on the 3D Quadrotor are similar to those on the 2D Quadrotor, again the generator mask converges the fastest but yields a final reward similar to that of the ray mask. The fraction of the relevant action space is on average $25\%$. Notably, on this environment the baseline does not learn a meaningful policy.
@@
Baseline Ray mask Generator mask Distributional mask
Seeker $-71.03\pm 19.670$ $-20.45\pm 16.39$ $18.60\pm 22.02$ $-13.66\pm 19.97$
2D Quadrotor $-0.80\pm 0.15$ $-0.26\pm 0.05$ $-0.25\pm 0.02$ $-0.23\pm 0.02$
3D Quadrotor $-1.00\pm 0.00$ $-0.63\pm 0.03$ $-0.68\pm 0.07$ $-0.66\pm 0.02$
4.3 Limitations

### v2

@@ 4.2 Generator mask
Zonotopes can be interpreted as the map of a hypercube in the generator space to a lower-dimensional space (see Section 3.2). The generator mask is based on exploiting this interpretation by letting the RL agent select actions in the hypercube of the generator space. Since the size of the output layer of the policy network cannot change during the learning process, we fix the dimension of the generator space. The generator mask requires the following assumption:
@@
The relevant action set $\mathcal{A}^{r}(s)$ is represented by a zonotope $\langle c(s),G(s)\rangle_{\mathcal{Z}}$, with $G\in\mathbb{R}^{N\times P}$ and $c\in\mathbb{R}^{N}$, and a state-invariant number of generators $P$.
Note that in practice, Assumption 3 can often be trivially fulfilled by choosing sufficiently many generators, and the number of generators $P$ is usually the output dimension of the parametrized policy. The domain of the policy $\pi_{\theta}(a|s)$ is the hypercube $\mathcal{A}^{l}=[-1,1]^{P}$, which can be interpreted as a latent action space, and the domain of the relevant policy $\pi_{\theta}^{r}(a^{r}|s)$ is a subset of the action space $\mathcal{A}^{r}\subseteq\mathcal{A}$ (see Figure 1(b)).
To derive the policy gradient of the generator mask, we assume:
@@
$\pi_{\theta}(a|s)$ is a parametrized normal distribution $\mathcal{N}(a;\mu_{\theta},\Sigma_{\theta})$.
@@
The relevant policy of the generator mask is
$\pi_{\theta}^{r}(a|s)=\mathcal{N}(a;G\mu_{\theta}+c,G\Sigma_{\theta}G^{T}).$ (11)
@@
The generator mask $g_{\mathrm{G}}:\mathcal{A}^{l}\rightarrow\mathcal{A}^{r}$ is:
$a^{r}=g_{\mathrm{G}}(a)=c+Ga,$ (12)
which is a linear function. Therefore, the proof directly follows from the linear transformation of multivariate normal distributions [Thm. 3.3.3][tong_multivariate_1990]. ∎
Note that the ray mask and generator mask are mathematically equivalent to the approach in [Krasowski2023b.ProvablySafeRLSurvey] if the relevant action set is constrained to intervals. Since $g_{\mathrm{G}}(a)$ is not bijective in general, we cannot derive the gradient through a change of variables, as for the ray mask.
@@
The policy gradient for $\pi_{\theta}^{r}(a^{r}|s)$ as defined in (11) with respect to $\mu_{\theta}$ and $\Sigma_{\theta}$ is
@@
The proposition is proven in Appendix A.2. ∎
Note that for the special case where $G^{-1}$ exists, i.e., $G$ is square and non-singular, the expressions in (13) and (14) simplify to $\nabla_{\mu_{\theta}}\log\pi_{\theta}(a|s)$, and $\nabla_{\Sigma_{\theta}}\log\pi_{\theta}(a|s)$, respectively, when Assumption 4 holds (see Proposition 5 in Appendix A.2). While the generator matrix will commonly not be square, as usually $P>N$, there are cases where $P=N$ is a valid choice, e.g., if there is a linear dependency between the action dimensions as for the 2D Quadrotor dynamics (see (40)).
@@ 4.3 Distributional mask
The intuition behind the distributional mask comes from discrete action masking, where the probability for irrelevant actions is set to 00 [Huang2022]. For continuous action spaces, we aim to achieve the same by ensuring that actions are only sampled from the relevant action set $\mathcal{A}^{r}$, by setting the density values of the relevant policy distribution $\pi_{\theta}^{r}(a^{r}|s)$ to zero for actions outside the relevant action set (see Figure 1(c)). For the one-dimensional case, this can be expressed by the truncated distribution [burkardt2018truncated]. In higher dimensions, the resulting policy distribution is
$\pi_{\theta}^{r}(a^{r}|s)=\frac{\phi(a,s)\pi_{\theta}(a|s)}{\int_{\mathcal{A}^{r}}\pi_{\theta}(\tilde{a}|s)\mathrm{d}\tilde{a}},$ (15)
@@
$\phi(a,s)=\begin{cases}1&\text{if }a\in\mathcal{A}^{r},\\ 0&\text{otherwise}.\end{cases}$ (16)
Since there is no closed form of this distribution, we employ Markov chain Monte Carlo sampling to sample actions from the policy. More specifically, we utilize the random direction hit-and-run algorithm: a geometric random walk that allows sampling from a non-negative, integrable function $f:\mathbb{R}^{N}\rightarrow\mathbb{R}_{+}$, while constraining the samples to a bounded set [zabinsky_hit-and-run_2013]. The algorithm iteratively chooses a random direction from the current point, computes the one-dimensional, truncated probability density of $f$ along this direction, and samples a new point from this density. The approach is particularly effective for high-dimensional spaces where other sampling methods might struggle with convergence or efficiency. For the distributional mask, $f$ is the policy $\pi_{\theta}(a|s)$, and $\mathcal{A}^{r}$ is the set. As suggested by [lovasz_hit-and-run_2006], we execute $N^{3}$ iterations before accepting the sample. To estimate the integral in (15), we use numerical integration with cubature [genz_adaptive_2003], which is a method to approximate the definite integral of a function $l:\mathbb{R}^{N}\rightarrow\mathbb{R}$ over a multidimensional geometric set.
@@
$\nabla_{\theta}\log\pi_{\theta}^{r}(a^{r}|s)=\nabla_{\theta}\log\pi_{\theta}(a|s)-\nabla_{\theta}\log\int_{\mathcal{A}^{r}}\pi_{\theta}(\tilde{a}|s)\mathrm{d}\tilde{a}.$ (17)
@@
Equation (17) is obtained by calculating the gradient of the logarithm of (15). The indicator function $\phi(a,s)$ is not continuous and differentiable, which would necessitate the use of the sub-gradient for learning. However, since $a^{r}\in\mathcal{A}^{r}$ always holds, the gradient has to be computed for the continuous part of $\pi_{\theta}^{r}(a^{r}|s)$ only, and thus $\phi(a,s)$ can be omitted. ∎
Since the gradient of the numeric integral $\int_{\mathcal{A}^{r}}\pi_{\theta}(\tilde{a}|s)\mathrm{d}\tilde{a}$ is intractable for zonotopes, we treat the integral as a constant in practice, and use $\nabla_{\theta}\log\pi_{\theta}^{r}(a^{r}|s)\approx\nabla_{\theta}\log\pi_{\theta}(a|s)$. We discuss potential improvements in Section 5.3.
@@ 5 Numerical experiments
We compare the three continuous action masking methods in four different environments: The simple and intuitive Seeker Reach-Avoid environment, the 2D Quadrotor environment to demonstrate the generalization from the continuous action masking approach in [Krasowski2023b.ProvablySafeRLSurvey], and the 3D Quadrotor and Mujoco Walker2D environment to show the applicability to action spaces of higher dimension. Because the derivation of the relevant action set is not trivial in practice, we selected four environments for which we could compute intuitive relevant action sets. In particular, for the Seeker and Quadrotor environments, the relevant action set is a safe action set since it is computed so that the agent does not collide (Seeker) or leave a control invariant set (2D and 3D Quadrotor). For the Walker2D environment, the relevant action set is state-independent and models a power supply constraint for the actuators.
@@ 5.1 Environments
We briefly introduce the three environments and their corresponding relevant action sets $\mathcal{A}^{r}$. Parameters and dynamics for the environments are detailed in Appendix A.5.
@@ 5.1.1 Seeker Reach-Avoid
This episodic environment features an agent navigating a 2D space, tasked with reaching a goal while avoiding a circular obstacle (see Figure 2). It is explicitly designed to provide an intuitive relevant action set $\mathcal{A}^{r}$. The system is defined as in Section 3.2: The dynamics for the position of the agent $s=\left[s_{x},s_{y}\right]^{T}$ and the action $a=\left[a_{x},a_{y}\right]^{T}$ is $\dot{s}=a$, and there are no disturbances.
$\mathcal{S}$$s^{*}$$\mathcal{O}$$\mathcal{S}_{\Delta t}$$s$$\mathcal{A}$$\mathcal{A}^{r}$
$\mathcal{A}$ action space
@@
Figure 2: The Seeker Reach-Avoid environment with state and action space. The agent (black) has to reach the goal (gray) while avoiding the obstacle (red). The center of the action space is illustrated by a cross and the relevant action set $\mathcal{A}^{r}$ for the current state is shown in green. The state set reachable at the next time step, by the relevant action set, is $\mathcal{S}_{\Delta t}$.
The environment is characterized by the position of the agent, the goal position $s^{*}$, the obstacle position $o$, and the obstacle radius $r_{o}$. These values are pseudo-randomly sampled at the beginning of each episode, with the constraints that the goal cannot be inside the obstacle and the obstacle blocks the direct path between the initial position of the agent and the goal. The reward for each time step is
$r(a,s)=\begin{cases}100&\text{if goal reached},\\ -100&\text{if collision occurred},\\ -1-\|s^{*}-s\|_{2}&\text{otherwise}.\end{cases}$ (18)
We compute the relevant action set $\mathcal{A}^{r}$ so that all actions that cause a collision with the obstacle or the boundary are excluded (see Appendix A.3.3).
@@
The 2D Quadrotor environment models a stabilization task and employs an action space where the two action dimensions are coupled, i.e., rotational movement is originating from differences between the action values and vertical movement is proportional to the sum of the action values. The relevant action set is computed based on the system dynamics and a relevant state set (see Appendix, Eq. (33)). The reward function is defined as
$r(a,s)=\exp\left(-\|s-s^{*}\|_{2}-\frac{0.01}{2}\left\|\left[\frac{a_{1}-a_{\mathrm{1,min}}}{a_{\mathrm{1,range}}},\frac{a_{2}-a_{\mathrm{2,min}}}{a_{\mathrm{2,range}}}\right]\right\|_{1}\right),$ (19)
where $s^{*}=\mathbf{0}$ is the stabilization goal state, $a=[a_{1},a_{2}]$ is the two-dimensional action, $a_{\mathrm{i,min}}$ is the lower bound for the actions in dimension $i$, and $a_{\mathrm{i,range}}$ is the absolute difference between the lower and upper bound for the actions in dimension $i$.
@@
The third environment models a stabilization task for a quadrotor defined in [Kaynama2015]. The quadrotor has four action dimensions. We use the same reward (see (19)) and the same calculation approach for the relevant action set (see Appendix, Eq. (33)) as for the 2D Quadrotor.
5.1.4 Mujoco Walker2D
The relevant action sets of the two Quadrotor and the Seeker environments are sets that only contain safe actions for the current state. Computing safe action sets requires considerable domain knowledge and for our case solving an optimization problems (see Appendix A.3). However, action masking is not restricted to safe relevant action sets, which we aim to demonstrate on the Mujoco Walker2D environment [todorov2012mujoco]. We extend the environment with a termination criterion, which ends an episode, when the the action violates the constraint $\|a\|_{2}\leq\alpha_{p}$. This can be viewed as constraining the cumulative power output on all joints to a maximum value $\alpha_{p}$. We motivate this constraint by having a battery with a maximum power output that should not be exceeded in practice. Accordingly, we define the relevant action set as the static set
$\mathcal{A}^{r}=\big{\{}a\big{|}\|a\|_{2}\leq\alpha_{p}\big{\}}.$ (20)
We under-approximate this relevant action set with a zonotope that consists of $36$ generators.
@@ 5.2 Results
The reported training results are based on ten random seeds for each configuration. Figure 3 shows the mean reward and the bootstrapped 95% confidence intervals [patterson2023empirical] for the four environments and Table 1 depicts the mean and standard deviation of the episode return during deployment. First, we present the results for the Seeker and Quadrotor environments for which the relevant action set is state-dependent and only includes safe actions. Then, we detail the results for the Walker2D environment with a static relevant action set.
For the Seeker and Quadotor environments, the baseline, i.e., PPO, converges significantly slower or not at all (see Figure 3). Additionally, the initial rewards when using action masking are significantly higher than for the baseline, indicating that the exploration is indeed constrained to relevant task-fulfilling actions. During deployment (see Table 1), the baseline episode returns are significantly lower than for action masking, while the three masking methods perform similarly. More specifically, for the Seeker environment, the relative volume of the relevant action set compared to the global action space is on average $70\%$, and all action masking methods converge significantly faster to high rewards than the baseline. The generator mask and the distributional mask achieve the highest final reward. Further, action replacement performs better than the baseline but significantly worse than masking. For the 2D Quadrotor, the relative volume is on average $28\%$, which is much smaller than for the Seeker environment. In the 2D Quadrotor environment, the ray mask, generator mask, and action replacement achieve the highest reward. While the baseline converges significantly slower, the final average reward is similar to the one of the distributional mask. Please note that the confidence interval for the baseline is significantly larger, because three of the ten runs do not converge and, on average, exhibit a reward of one throughout training. Additionally, we observed that if we constrain the relevant action set to intervals for this environment, our optimization problem in (33) often renders infeasible, because the maximal relevant action set cannot be well approximated by an interval. Thus, the masking method proposed in [Krasowski2023b.ProvablySafeRLSurvey] is not suitable for the 2D Quadrotor task. The results on the 3D Quadrotor are similar to those on the 2D Quadrotor; again, the generator mask converges the fastest but yields a final reward similar to that of the ray mask and action replacement. The relative volume of the relevant action set to the global action space is on average $25\%$. Notably, in this environment, the baseline does not learn a meaningful policy. Based on these three environments with state-dependent relevant action sets, it seems that action masking performs better than action replacement when the relative volume is not too small.
@@
Seeker 2D Quad. 3D Quad. Walker2D
Baseline $-71.03\pm 19.67$ $-0.80\pm 0.15$ $-1.00\pm 0.00$ $1.19\pm 0.06$
Replacement $-60.21\pm 19.98$ $-0.25\pm 0.03$ $-0.68\pm 0.09$ $234.93\pm 129.5$
Ray $-20.45\pm 16.39$ $-0.26\pm 0.05$ $-0.63\pm 0.03$ $1941.82\pm 992.83$
Generator $18.60\pm 22.02$ $-0.25\pm 0.02$ $-0.68\pm 0.07$ $1443.62\pm 702.7$
Distributional $-13.66\pm 19.97$ $-0.23\pm 0.02$ $-0.66\pm 0.02$
The training results of the Walker2D experiment are shown in Figure 3. While the generator and ray mask both learn a performant policy, the ray mask outperforms by a significant margin. The lower performance of the generator mask is likely due to the high-dimensional generator space with $36$ dimensions. This is supported by initial experiments with $12$ generators (i.e., a more conservative under-approximation of the $L_{2}$-norm) where the generator mask performed better compared to the ray mask. Replacement performs significantly worse than the two masking approaches, and the PPO baseline does not learn a meaningful policy since the environment is frequently reset due to violations of the power constraint in (20). The frequent terminations occur since in six action dimensions the relative volume of the unit ball compared to the unit box is $\approx 8\%$, i.e., more than $92\%$ of actions are violating the power constraint. To compare masking to a learning baseline, we also evaluated standard PPO without constraints, which performs slightly better than ray masking but almost always uses actions outside $\mathcal{A}^{r}$ (see Appendix A.7). The deployment results in Table 1 reflect similar results as the training; the ray mask achieves better rewards than the generator mask, followed by replacement, and the baseline performs the worst. We excluded the distributional mask for the Walker2D, since its computation time is approximately $170$ times slower than the baseline, compared to a $1.6$ increase for the generator, $2.7$ for the ray mask, and $2.5$ for action replacement (see Table 3). The severely increased computational cost for the distributional mask arises from the geometric random walks, which scale cubically with action dimensions.
5.3 Discussion and limitations